OpenStudy (anonymous):

Which of the following is the equation of the line passing through the points C(-1, 1) and D(7 -1). Geometry help!

OpenStudy (anonymous):

@GirlByte

OpenStudy (anonymous):

A- y+1=1/4(x-1) B- y-1=-1/4(x+1) C- y-1=4(x+1) D- y-1=1/4(x+1)